OpenStudy (anonymous):

solve each equation on the interval (0,2pie) use exact values where possible or give solutions correct to four decimal places 4sin^2x=1

OpenStudy (anonymous):

start with \[\sin^2(x)=\frac{1}{4}\] and then solve \[sin(x)=\pm\frac{1}{2}\]

OpenStudy (anonymous):

what about cos2x-sinx=1

hero (hero):

I will post the solution in a sec

hero (hero):

Rohangrr, please allow me to post the complete solution

OpenStudy (anonymous):

and perhaps even a correct one

hero (hero):

It will be correct smartie

OpenStudy (anonymous):

okay! @Hero

OpenStudy (anonymous):

lol i have more questions

hero (hero):

@satellite73 , I know I make mistakes, but I usually correct them.

OpenStudy (anonymous):

i am sure it will be, i was referring to the fact that the last two rohangrr answers were random i am sure you (hero) will have no trouble with this

hero (hero):

\[\cos^2x - \sin x = 1\]\[(1 - \sin^2x) - \sin x = 1\]\[1 - \sin^2 x - \sin x = 1\]\[-\sin^2x - \sin x = 1 - 1\]\[-\sin^2x - \sin x = 0\]\[-(\sin^2x + \sin x) = 0\]\[\sin^2x + \sin x = 0\]\[\sin x(\sin x + 1) = 0\] Now use zero product property to continue solving for x

hero (hero):

sin x = 0 sin x + 1 = 0

OpenStudy (anonymous):

thank u